# Spectral selectors on lens spaces and applications to the geometry of the group of contactomorphisms

### v1

@@ 4. Non-shortening of the standard Reeb flow with respect to the discriminant and oscillation norms
Recall from [9] that the discriminant norm $\nu_{\operatorname{dis}}$ on the universal cover $\widetilde{\operatorname{Cont}_{0}}(M,\xi)$ of the identity component of the contactomorphism group of a closed contact manifold $(M,\xi)$ is the word norm associated to the generating set $\mathcal{D}\subset\widetilde{\operatorname{Cont}_{0}}(M,\xi)$ formed by elements $\widetilde{\phi}$ that can be represented by an embedded contact isotopy, i.e. a contact isotopy $\{\phi_{t}\}_{t\in[0,1]}$ such that $\phi_{t}\circ\phi_{s}^{-1}$ has no discriminant points for all $s\neq t\in[0,1]$. Recall also from [5] that the discriminant length of a contact isotopy $\{\phi_{t}\}_{t\in[0,1]}$ is the minimal $N$ such that there is a decomposition $10=t_{0}<\cdots<t_{N}=$ 1 of the time interval $[0,1]$ with $\{\phi_{t}\}_{t\in[t_{j},t_{j+1}]}$ embedded for all $j=0,\cdots,N-1$.
Consider the discriminant norm on $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$. For every positive real number $T$ we have
$\nu_{\operatorname{dis}}(\widetilde{r_{T}})\geq\frac{2n\lceil\frac{T}{2\pi}\rceil+1}{2n+1}\,.$
@@
$2n\,\left\lceil\frac{T}{2\pi}\right\rceil=\mu(\widetilde{r_{T}})\leq\sum_{j=1}^{N}\mu(\widetilde{\phi_{j}})+N-1\leq 2nN+N-1\,.$ (17)
@@
$\{r_{Tt}\}_{t\in[0,1]}=\{\underbrace{r_{T_{0}t}\circ\cdots\circ r_{T_{0}t}}_{\left\lfloor\frac{k}{2\pi}\,T\right\rfloor+1}\}_{t\in[0,1]}$ (19)
and so, since $\{r_{T_{0}t}\}_{t\in[0,1]}\in\mathcal{D}$, $\{r_{Tt}\}_{t\in[0,1]}$ has discriminant length smaller or equal than $\left\lfloor\frac{k}{2\pi}\,T\right\rfloor+1$; this length is actually equal to $\left\lfloor\frac{k}{2\pi}\,T\right\rfloor+1$, because for any interval $[t_{0},t_{1}]$ of length $t_{1}-t_{0}\geq\frac{2\pi}{k}$ the contact isotopy $\{r_{t}\}_{t\in[t_{0},t_{1}]}$ is not embedded. For instance, in the case of projective space the discriminant length of $\{r_{2\pi mt}\}_{t\in[0,1]}$ is thus $14m+$ 1, while (18) only gives $\nu_{\operatorname{dis}}(\widetilde{r_{2\pi m}})\geq m$. In this section we prove the optimal estimates for the discriminant and oscillation lengths of the Reeb flow of lens spaces using the spectral selectors defined in Section 3. The main advantage of using the spectral selectors is that while (by Section 2 (v)) the non-linear Maslov index only jumps in the presence of discriminant points of the lift of a contact isotopy of $(L_{k}^{2n-1},\xi_{0})$ to the sphere, so that in particular for instance $\mu(\widetilde{r_{T}})=2n\,\left\lceil\frac{T}{2\pi}\right\rceil$, the spectral selectors allow to distinguish $\widetilde{r_{T}}$ also for different values of $T$ in $[0,2\pi]$, indeed by Theorem 1.1 (v) we have for instance $c_{0}(\widetilde{r_{T}})=T$.
@@
For any element $\widetilde{\phi}$ of $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$, if $\widetilde{\phi}\in\mathcal{D}$ then $c_{0}(\widetilde{\phi})<\frac{2\pi}{k}$.
@@
If $\widetilde{\phi}\in\mathcal{D}$ then $\widetilde{\phi}$ can be represented by a contact isotopy $\{\phi_{t}\}_{t\in[0,1]}$ such that $\phi_{t}$ does not have discriminant points for all $t\in(0,1]$. Suppose by contradiction that $c_{0}(\widetilde{\phi})\geq\frac{2\pi}{k}$, and for $s\in[0,1]$ let $\widetilde{\phi}_{s}=[\{\phi_{st}\}_{t\in[0,1]}]$. Since $c_{0}(\widetilde{\phi}_{1})=c_{0}(\widetilde{\phi})\geq\frac{2\pi}{k}$ and, by Theorem 1.1 (ii), $c_{0}(\widetilde{\phi}_{0})=c_{0}(\widetilde{\mathrm{id}})=0$, by continuity of $c_{0}$ (Theorem 1.1 (viii)) there is a value of $s$ in $(0,1]$ such that $c_{0}(\widetilde{\phi}_{s})=\frac{2\pi}{k}$. But then, by spectrality (Theorem 1.1 (i)), $\frac{2\pi}{k}$ belongs to $\bar{\mathcal{A}}(\widetilde{\phi}_{s})$. This means that $\phi_{s}$ has discriminant points, which is a contradiction. ∎
We have seen above that $\{r_{Tt}\}_{t\in[0,1]}$ has discriminant length $\left\lfloor\frac{k}{2\pi}\,T\right\rfloor+1$. In order to prove that it is a geodesic we thus have to show that
@@
Let $\nu_{\operatorname{dis}}(\widetilde{r_{T}})=N$, and write $\widetilde{r_{T}}=\prod_{j=1}^{N}\widetilde{\phi_{j}}$ with $\widetilde{\phi_{j}}\in\mathcal{D}$ for all $j$. By Theorem 1.1 (v), (ix) and Lemma 4 we then have
@@
We now show that $\{r_{Tt}\}_{t\in[0,1]}$ is a geodesic with respect to the oscillation norm. Recall from [9] that the oscillation pseudonorm $\nu_{\operatorname{osc}}$ on the universal cover $\widetilde{\operatorname{Cont}_{0}}(M,\xi)$ of the identity component of the contactomorphism group of a closed contact manifold $(M,\xi)$ is defined as follows. Let $\mathcal{D}_{+}$ and $\mathcal{D}_{-}$ be the sets of elements of $\widetilde{\operatorname{Cont}_{0}}(M,\xi)$ that can be represented respectively by an embedded non-negative or non-positive contact isotopy. It is proved in [9] that every element $\widetilde{\phi}$ of $\widetilde{\operatorname{Cont}_{0}}(M,\xi)$ can be written as $\widetilde{\phi}=\prod_{j=1}^{N}\widetilde{\phi_{j}}$ with $\widetilde{\phi_{j}}\in\mathcal{D}_{+}$ or $\widetilde{\phi_{j}}\in\mathcal{D}_{-}$ for every $j$. We denote by $\nu_{+}(\widetilde{\phi})$ and $\nu_{-}(\widetilde{\phi})$ respectively the minimal number of elements of $\mathcal{D}_{+}$ and minus the minimal number of elements of $\mathcal{D}_{-}$ in such a decomposition. The oscillation pseudonorm is then defined by
@@
for $\widetilde{\phi}\neq\widetilde{\mathrm{id}}$, and $\nu_{\operatorname{osc}}(\widetilde{\mathrm{id}})=0$. By [9, Proposition 3.2], the oscillation pseudonorm on $\widetilde{\operatorname{Cont}_{0}}(M,\xi)$ is non-degenerate if and only if $(M,\xi)$ is orderable; it is thus a norm for lens spaces. Recall also from [5] that the oscillation length of a contact isotopy $\{\phi_{t}\}_{t\in[0,1]}$ is the sum of +($\mathcal{L}_{+}\big{(}\{\phi_{t}\}_{t\in[0,1]}\big{)}$ and ($\mathcal{L}_{-}\big{(}\{\phi_{t}\}_{t\in[0,1]}\big{)}$, where +($\mathcal{L}_{+}\big{(}\{\phi_{t}\}_{t\in[0,1]}\big{)}$ is the minimal $N_{+}$ for which there is $N\geq N_{+}$ and a decomposition $10=t_{0}<\cdots<t_{N}=$ 1 with each $\{\phi_{t}\}_{t\in[t_{j},t_{j+1}]}$ embedded and non-negative or non-positive and exactly $N_{+}$ of them non-negative, and ($\mathcal{L}_{-}\big{(}\{\phi_{t}\}_{t\in[0,1]}\big{)}$ is the minimal $N_{-}$ for which there is $N\geq N_{-}$ and a decomposition $10=t_{0}<\cdots<t_{N}=$ 1 with each $\{\phi_{t}\}_{t\in[t_{j},t_{j+1}]}$ embedded and non-negative or non-positive and exactly $N_{-}$ of them non-positive.
Consider now as above the Reeb flow $\widetilde{r_{T}}=[\{r_{Tt}\}_{t\in[0,1]}]$ on $(L_{k}^{2n-1},\xi_{0})$. The decomposition (19) shows that $\nu_{-}(\widetilde{r_{T}})=0$ and $\nu_{+}(\widetilde{r_{T}})\leq\left\lfloor\frac{k}{2\pi}\,T\right\rfloor+1$, thus
$\nu_{\operatorname{osc}}(\widetilde{r_{T}})=\nu_{+}(\widetilde{r_{T}})\leq\left\lfloor\frac{k}{2\pi}\,T\right\rfloor+1\,.$
In order to show that $\{r_{Tt}\}_{t\in[0,1]}$ is a geodesic with respect to the oscillation norm we thus have to show that
@@
Let $\nu_{+}(\widetilde{r_{T}})=N_{+}$, and write $\widetilde{r_{T}}=\prod\limits_{j=1}^{N}\widetilde{\phi}_{j}$ with $\widetilde{\phi}_{j}\in\mathcal{D}_{\pm}$ for all $j$ and with exactly $N_{+}$ of the $\widetilde{\phi_{j}}$ in $\mathcal{D}_{+}$. Denote such elements by $\widetilde{\phi_{\sigma(1)}}$, $\cdots$, $\widetilde{\phi_{\sigma(N_{+})}}$. Then $\widetilde{r_{T}}\leq\prod\limits_{j=1}^{N_{+}}\widetilde{\phi}_{\sigma(j)}$, and so by Theorem 1.1 (v), (vii), (ix) and Lemma 4 we have
@@ 5. A spectral pseudonorm
Let $c_{-}=c_{-2n+1}$ and $c_{+}=c_{0}$, and define $\nu:\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})\rightarrow\frac{2\pi}{k}\cdot\mathbb{Z}$ by
$\nu(\widetilde{\phi})=\max\left\{\,\left\lceil c_{+}(\widetilde{\phi})\right\rceil_{\frac{2\pi}{k}}\,,\,-\left\lfloor c_{-}(\widetilde{\phi})\right\rfloor_{\frac{2\pi}{k}}\,\right\}\,.$
@@
The map $\nu:\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})\rightarrow\frac{2\pi}{k}\cdot\mathbb{Z}$ is a stably unbounded conjugation invariant pseudonorm compatible with the partial order $\leq$.
@@
We first show that for every $\widetilde{\phi}$ we have $\nu(\widetilde{\phi})\geq 0$. Suppose by contradiction that $\nu(\widetilde{\phi})<0$. Then $\left\lceil c_{+}(\widetilde{\phi})\right\rceil_{\frac{2\pi}{k}}<0$, thus $c_{+}(\widetilde{\phi})<0$, and $-\left\lfloor c_{-}(\widetilde{\phi})\right\rfloor_{\frac{2\pi}{k}}<0$, thus $c_{-}(\widetilde{\phi})>0$. But this contradicts the fact that, since the sequence $c_{j}$ is non-decreasing, $c_{-}(\widetilde{\phi})\leq c_{+}(\widetilde{\phi})$. The triangle inequality $\nu(\widetilde{\phi}\cdot\widetilde{\psi})\leq\nu(\widetilde{\phi})+\nu(\widetilde{\psi})$ follows from Theorem 1.1 (ix) and (xi), and symmetry, i.e. $\nu(\widetilde{\phi})=\nu(\widetilde{\phi}^{-1})$, from Theorem 1.1 (xi). This shows that $\nu$ is a pseudonorm. Invariance by conjugation follows from Theorem 1.1 (x) and (xi). The pseudonorm $\nu$ is stably unbounded, indeed Theorem 1.1 (v) implies that
$\nu\Big{(}{\widetilde{r_{\frac{2\pi}{k}}}}^{m}\Big{)}=\nu\big{(}\widetilde{r_{m\frac{2\pi}{k}}}\big{)}=m$
for every positive integer $m$, thus posing $\sigma=\widetilde{r_{\frac{2\pi}{k}}}$ we have $\lim_{m\to\infty}\frac{\nu(\sigma^{m})}{m}=1\neq 0$. Finally, the fact that $\nu$ is compatible with the partial order $\leq$ follows from Theorem 1.1 (vii) and (xi). ∎
@@
For every element $\widetilde{\phi}$ of $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$ we have
$\nu(\widetilde{\phi})\leq\frac{2\pi}{k}\cdot\nu_{\operatorname{osc}}(\widetilde{\phi})\,.$
@@
$c_{+}(\widetilde{\phi})\leq c_{+}\Big{(}\prod_{j=1}^{N_{+}}\widetilde{\phi}_{\sigma(j)}\Big{)}\leq\sum\limits_{j=1}^{N_{+}}\left\lceil c_{+}(\widetilde{\phi}_{\sigma(j)})\right\rceil_{\frac{2\pi}{k}}\leq\frac{2\pi}{k}\cdot N_{+}\,.$
Similarly, setting $\nu_{-}(\widetilde{\phi})=-N_{-}$ we have $c_{+}(\widetilde{\phi}^{-1})\leq\frac{2\pi}{k}\cdot N_{-}$ and so, by Theorem 1.1 (xi),
$-\left\lfloor c_{-}(\widetilde{\phi})\right\rfloor_{\frac{2\pi}{k}}=\left\lceil c_{+}\left(\widetilde{\phi}^{-1}\right)\right\rceil_{\frac{2\pi}{k}}\leq\frac{2\pi}{k}\cdot N_{-}\,.$
@@
$\nu(\widetilde{\phi})\leq\frac{2\pi}{k}\,\max\Big{\{}\nu_{+}(\widetilde{\phi})\,,\,-\nu_{-}(\widetilde{\phi})\Big{\}}\leq\frac{2\pi}{k}\big{(}\nu_{+}(\widetilde{\phi})-\nu_{-}(\widetilde{\phi})\big{)}=\frac{2\pi}{k}\cdot\nu_{\operatorname{osc}}(\widetilde{\phi})\,.$
We do not know whether the pseudonorm $\nu$ is non-degenerate, i.e. whether $\nu(\widetilde{\phi})=0$ if and only if $\widetilde{\phi}=\widetilde{\mathrm{id}}$. Indeed, by the definition of $\nu$ we have that $\nu(\widetilde{\phi})=0$ if and only if $c_{+}(\widetilde{\phi})=c_{-}(\widetilde{\phi})=0$, which by Theorem 1.1 (iv) only implies that $\Pi(\widetilde{\phi})$ is the identity. On the other hand, the induced conjugation invariant pseudonorm on $\operatorname{Cont}_{0}(L_{k}^{2n-1},\xi_{0})$, i.e. the pseudonorm $\nu_{\ast}$ defined by
@@
For every $\phi\in\operatorname{Cont}_{0}(L_{k}^{2n-1},\xi_{0})$ we have $\nu_{\ast}(\phi)\leq 2\pi+\frac{2\pi}{k}$.

### v2

@@ 4. Non-shortening of the standard Reeb flow with respect to the discriminant and oscillation norms
Recall from [10] that the discriminant norm $\nu_{\operatorname{dis}}$ on the universal cover $\widetilde{\operatorname{Cont}_{0}}(M,\xi)$ of the identity component of the contactomorphism group of a closed contact manifold $(M,\xi)$ is the word norm associated to the generating set $\mathcal{D}$ formed by elements $\widetilde{\phi}$ that can be represented by an embedded contact isotopy, i.e. a contact isotopy $\{\phi_{t}\}_{t\in[0,1]}$ such that $\phi_{t}\circ\phi_{s}^{-1}$ has no discriminant points for all $s\neq t\in[0,1]$. Recall also from [6] that the discriminant length of a contact isotopy $\{\phi_{t}\}_{t\in[0,1]}$ is the minimal $N$ such that there is a decomposition $10=t_{0}<\cdots<t_{N}=$ 1 of the time interval $[0,1]$ with $\{\phi_{t}\}_{t\in[t_{j},t_{j+1}]}$ embedded for all $j=0,\cdots,N-1$.
Consider the discriminant norm on $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$. For every positive real number $T$ we have
$\nu_{\operatorname{dis}}(\widetilde{r_{T}})\geq\frac{2n\lceil\frac{T}{2\pi}\rceil+1}{2n+1}\,.$ (17)
@@
$2n\,\left\lceil\frac{T}{2\pi}\right\rceil=\mu(\widetilde{r_{T}})\leq\sum_{j=1}^{N}\mu(\widetilde{\phi_{j}})+N-1\leq 2nN+N-1\,.$
@@
$\{r_{Tt}\}_{t\in[0,1]}=\{\underbrace{r_{T_{0}t}\circ\cdots\circ r_{T_{0}t}}_{\left\lfloor\frac{k}{2\pi}\,T\right\rfloor+1}\}_{t\in[0,1]}\,.$ (19)
Since the minimal period of a closed Reeb orbit of $\alpha_{0}$ on $L_{k}^{2n-1}(\underline{w})$ is $\frac{2\pi}{k}$ we have that $\{r_{T_{0}t}\}_{t\in[0,1]}\in\mathcal{D}$, and so $\{r_{Tt}\}_{t\in[0,1]}$ has discriminant length smaller or equal than $\left\lfloor\frac{k}{2\pi}\,T\right\rfloor+1$; this length is actually equal to $\left\lfloor\frac{k}{2\pi}\,T\right\rfloor+1$, because for any interval $[t_{0},t_{1}]$ of length $t_{1}-t_{0}\geq\frac{2\pi}{k}$ the contact isotopy $\{r_{t}\}_{t\in[t_{0},t_{1}]}$ is not embedded. For instance, in the case of projective space the discriminant length of $\{r_{2\pi mt}\}_{t\in[0,1]}$ is thus $12m+$ 1, while (18) only gives $\nu_{\operatorname{dis}}(\widetilde{r_{2\pi m}})\geq m$. In this section we prove the optimal estimates for the discriminant and oscillation lengths of the standard Reeb flow of lens spaces with equal weights using the spectral selectors defined in Section 3. The main advantage of using the spectral selectors is that while (by Section 2 (v)) the non-linear Maslov index only jumps in the presence of discriminant points of the lift of a contact isotopy of a lens space to the sphere, so that in particular for instance $\mu(\widetilde{r_{T}})=2n\,\left\lceil\frac{T}{2\pi}\right\rceil$, the spectral selectors allow to distinguish $\widetilde{r_{T}}$ also for different values of $T$ in $[0,2\pi]$, indeed by Theorem 1.1 (v) we have for instance $c_{0}(\widetilde{r_{T}})=T$.
@@
For any element $\widetilde{\phi}$ of $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$, if $\widetilde{\phi}\in\mathcal{D}$ then $c_{0}(\widetilde{\phi})<T_{\underline{w}}$.
@@
If $\widetilde{\phi}\in\mathcal{D}$ then $\widetilde{\phi}$ can be represented by a contact isotopy $\{\phi_{t}\}_{t\in[0,1]}$ such that $\phi_{t}$ does not have discriminant points for all $t\in(0,1]$. Suppose by contradiction that $c_{0}(\widetilde{\phi})\geq T_{\underline{w}}$, and for $s\in[0,1]$ let $\widetilde{\phi}_{s}=[\{\phi_{st}\}_{t\in[0,1]}]$. Since $c_{0}(\widetilde{\phi}_{1})=c_{0}(\widetilde{\phi})\geq T_{\underline{w}}$ and, by Theorem 1.1 (ii), $c_{0}(\widetilde{\phi}_{0})=c_{0}(\widetilde{\mathrm{id}})=0$, by continuity of $c_{0}$ (Theorem 1.1 (viii)) there is a value of $s$ in $(0,1]$ such that $c_{0}(\widetilde{\phi}_{s})=T_{\underline{w}}$. But then, by spectrality (Theorem 1.1 (i)), $T_{\underline{w}}$ belongs to $\bar{\mathcal{A}}(\widetilde{\phi}_{s})$. This means that $\phi_{s}$ has discriminant points, which is a contradiction. ∎
We can now prove that for every real number $T$ the Reeb flow $\{r_{Tt}\}_{t\in[0,1]}$ of the standard contact form $\alpha_{0}$ on a lens space of the form $L_{k}^{2n-1}(w,\cdots,w)$ is a geodesic for the discriminant norm. We have seen above that $\{r_{Tt}\}_{t\in[0,1]}$ has discriminant length $\left\lfloor\frac{k}{2\pi}\,T\right\rfloor+1$. In order to prove that it is a geodesic we thus have to show that
@@
Let $\nu_{\operatorname{dis}}(\widetilde{r_{T}})=N$, and write $\widetilde{r_{T}}=\prod_{j=1}^{N}\widetilde{\phi_{j}}$ with $\widetilde{\phi_{j}}\in\mathcal{D}$ for all $j$. By Theorem 1.1 (v), (ix) and Lemma 4, and since $T_{\underline{w}}=\frac{2\pi}{k}$ for $\underline{w}=(w,\cdots,w)$, we then have
@@
We now show that the standard Reeb flow $\{r_{Tt}\}_{t\in[0,1]}$ on a lens space of the form $L_{k}^{2n-1}(w,\cdots,w)$ is a geodesic with respect to the oscillation norm. Recall from [10] that the oscillation pseudonorm $\nu_{\operatorname{osc}}$ on the universal cover $\widetilde{\operatorname{Cont}_{0}}(M,\xi)$ of the identity component of the contactomorphism group of a closed contact manifold $(M,\xi)$ is defined as follows. Let $\mathcal{D}_{+}$ and $\mathcal{D}_{-}$ be the sets of elements of $\widetilde{\operatorname{Cont}_{0}}(M,\xi)$ that can be represented respectively by an embedded non-negative or non-positive contact isotopy. It is proved in [10] that every element $\widetilde{\phi}$ of $\widetilde{\operatorname{Cont}_{0}}(M,\xi)$ can be written as $\widetilde{\phi}=\prod_{j=1}^{N}\widetilde{\phi_{j}}$ with $\widetilde{\phi_{j}}\in\mathcal{D}_{+}$ or $\widetilde{\phi_{j}}\in\mathcal{D}_{-}$ for every $j$. We denote by $\nu_{+}(\widetilde{\phi})$ and $\nu_{-}(\widetilde{\phi})$ respectively the minimal number of elements of $\mathcal{D}_{+}$ and minus the minimal number of elements of $\mathcal{D}_{-}$ in such a decomposition. The oscillation pseudonorm is then defined by
@@
for $\widetilde{\phi}\neq\widetilde{\mathrm{id}}$, and $\nu_{\operatorname{osc}}(\widetilde{\mathrm{id}})=0$. By [10, Proposition 3.2], the oscillation pseudonorm on $\widetilde{\operatorname{Cont}_{0}}(M,\xi)$ is non-degenerate if and only if $(M,\xi)$ is orderable; it is thus a norm for lens spaces. Recall also from [6] that the oscillation length of a contact isotopy $\{\phi_{t}\}_{t\in[0,1]}$ is the sum of +($\mathcal{L}_{+}\big{(}\{\phi_{t}\}_{t\in[0,1]}\big{)}$ and ($\mathcal{L}_{-}\big{(}\{\phi_{t}\}_{t\in[0,1]}\big{)}$, where +($\mathcal{L}_{+}\big{(}\{\phi_{t}\}_{t\in[0,1]}\big{)}$ is the minimal $N_{+}$ for which there is $N\geq N_{+}$ and a decomposition $10=t_{0}<\cdots<t_{N}=$ 1 with each $\{\phi_{t}\}_{t\in[t_{j},t_{j+1}]}$ embedded and non-negative or non-positive and exactly $N_{+}$ of them non-negative, and ($\mathcal{L}_{-}\big{(}\{\phi_{t}\}_{t\in[0,1]}\big{)}$ is the minimal $N_{-}$ for which there is $N\geq N_{-}$ and a decomposition $10=t_{0}<\cdots<t_{N}=$ 1 with each $\{\phi_{t}\}_{t\in[t_{j},t_{j+1}]}$ embedded and non-negative or non-positive and exactly $N_{-}$ of them non-positive.
Consider now as above the class $\widetilde{r_{T}}=[\{r_{Tt}\}_{t\in[0,1]}]$ of the Reeb flow of $\alpha_{0}$ on a lens space $\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$. Similarly as before, the oscillation length of $\{r_{Tt}\}_{t\in[0,1]}$ is $\left\lfloor\frac{k}{2\pi}\,T\right\rfloor+1$. The decomposition (19) shows moreover that $\nu_{-}(\widetilde{r_{T}})=0$, and thus $\nu_{\operatorname{osc}}(\widetilde{r_{T}})=\nu_{+}(\widetilde{r_{T}})$. In order to show that the Reeb flow $\{r_{Tt}\}_{t\in[0,1]}$ of $\alpha_{0}$ on a lens space of the form $L_{k}^{2n-1}(w,\cdots,w)$ is a geodesic with respect to the oscillation norm we thus have to show that
@@
Let $\nu_{+}(\widetilde{r_{T}})=N_{+}$, and write $\widetilde{r_{T}}=\prod\limits_{j=1}^{N}\widetilde{\phi}_{j}$ with $\widetilde{\phi}_{j}\in\mathcal{D}_{\pm}$ for all $j$ and with exactly $N_{+}$ of the $\widetilde{\phi_{j}}$ in $\mathcal{D}_{+}$. Denote such elements by $\widetilde{\phi_{\sigma(1)}}$, $\cdots$, $\widetilde{\phi_{\sigma(N_{+})}}$. Then $\widetilde{r_{T}}\leq\prod\limits_{j=1}^{N_{+}}\widetilde{\phi}_{\sigma(j)}$, and so by Theorem 1.1 (v), (vii), (ix) and Lemma 4, and since $T_{\underline{w}}=\frac{2\pi}{k}$ for $\underline{w}=(w,\cdots,w)$, we have
@@
[…]
For a general lens space $\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ the above discussion implies that the discriminant and oscillation norms of $\widetilde{r_{T}}=[\{r_{Tt}\}_{t\in[0,1]}]$ are greater or equal than $\left\lfloor\frac{T}{T_{\underline{w}}}\right\rfloor+1$, while the discriminant and oscillation lengths of $\{r_{Tt}\}_{t\in[0,1]}$ are equal to $\left\lfloor\frac{k}{2\pi}\,T\right\rfloor+1$. We do not know thus if the standard Reeb flow on general lens spaces is a geodesic for the discriminant and oscillation norms. The gap for general weights between the minimal period $\frac{2\pi}{k}$ of a closed Reeb orbit and the period $T_{\underline{w}}$ of the Reeb flow seems to suggest that it might be possible to shorten the Reeb flow. It would be interesting to investigate if this is indeed the case, or if other methods could be used to prove that the Reeb flow is still a geodesic.
@@ 5. A spectral pseudonorm
Let $c_{-}=c_{-2n+1}$ and $c_{+}=c_{0}$, and define $\nu:\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}\rightarrow T_{\underline{w}}\cdot\mathbb{Z}$ by
$\nu(\widetilde{\phi})=\max\left\{\,\left\lceil c_{+}(\widetilde{\phi})\right\rceil_{T_{\underline{w}}}\,,\,-\left\lfloor c_{-}(\widetilde{\phi})\right\rfloor_{T_{\underline{w}}}\,\right\}\,.$
@@
The map $\nu:\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}\rightarrow T_{\underline{w}}\cdot\mathbb{Z}$ is a stably unbounded conjugation invariant pseudonorm compatible with the partial order $\leq$.
@@
We first show that for every $\widetilde{\phi}$ we have $\nu(\widetilde{\phi})\geq 0$. Suppose by contradiction that $\nu(\widetilde{\phi})<0$. Then $\left\lceil c_{+}(\widetilde{\phi})\right\rceil_{T_{\underline{w}}}<0$, thus $c_{+}(\widetilde{\phi})<0$, and $-\left\lfloor c_{-}(\widetilde{\phi})\right\rfloor_{T_{\underline{w}}}<0$, thus $c_{-}(\widetilde{\phi})>0$. But this contradicts the fact that, since the sequence $c_{j}$ is non-decreasing, $c_{-}(\widetilde{\phi})\leq c_{+}(\widetilde{\phi})$. The triangle inequality $\nu(\widetilde{\phi}\cdot\widetilde{\psi})\leq\nu(\widetilde{\phi})+\nu(\widetilde{\psi})$ follows from Theorem 1.1 (ix) and (xi), and symmetry $\nu(\widetilde{\phi})=\nu(\widetilde{\phi}^{-1})$ from Theorem 1.1 (xi). This shows that $\nu$ is a pseudonorm. Invariance by conjugation follows from Theorem 1.1 (x) and (xi). The pseudonorm $\nu$ is stably unbounded, indeed Theorem 1.1 (v) implies that
$\nu\Big{(}{\widetilde{r_{T_{\underline{w}}}}}^{m}\Big{)}=\nu\big{(}\widetilde{r_{mT_{\underline{w}}}}\big{)}=m\,T_{\underline{w}}$
for every positive integer $m$, thus posing $\sigma=\widetilde{r_{T_{\underline{w}}}}$ we have $\lim_{m\to\infty}\frac{\nu(\sigma^{m})}{m}=T_{\underline{w}}\neq 0$. Finally, the fact that $\nu$ is compatible with the partial order $\leq$ follows from Theorem 1.1 (vii) and (xi). ∎
@@
For every element $\widetilde{\phi}$ of $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ we have
$\nu(\widetilde{\phi})\leq T_{\underline{w}}\cdot\nu_{\operatorname{osc}}(\widetilde{\phi})\,.$
@@
$c_{+}(\widetilde{\phi})\leq c_{+}\Big{(}\prod_{j=1}^{N_{+}}\widetilde{\phi}_{\sigma(j)}\Big{)}\leq\sum\limits_{j=1}^{N_{+}}\left\lceil c_{+}(\widetilde{\phi}_{\sigma(j)})\right\rceil_{T_{\underline{w}}}\leq T_{\underline{w}}\cdot N_{+}\,.$
Similarly, setting $\nu_{-}(\widetilde{\phi})=-N_{-}$ we have $c_{+}(\widetilde{\phi}^{-1})\leq T_{\underline{w}}\cdot N_{-}$ and so, by Theorem 1.1 (xi),
$-\left\lfloor c_{-}(\widetilde{\phi})\right\rfloor_{T_{\underline{w}}}=\left\lceil c_{+}\left(\widetilde{\phi}^{-1}\right)\right\rceil_{T_{\underline{w}}}\leq T_{\underline{w}}\cdot N_{-}\,.$
@@
$\nu(\widetilde{\phi})\leq T_{\underline{w}}\,\max\Big{\{}\nu_{+}(\widetilde{\phi})\,,\,-\nu_{-}(\widetilde{\phi})\Big{\}}\leq T_{\underline{w}}\big{(}\nu_{+}(\widetilde{\phi})-\nu_{-}(\widetilde{\phi})\big{)}=T_{\underline{w}}\cdot\nu_{\operatorname{osc}}(\widetilde{\phi})\,.$
We do not know whether the pseudonorm $\nu$ is non-degenerate, i.e. whether $\nu(\widetilde{\phi})=0$ if and only if $\widetilde{\phi}=\widetilde{\mathrm{id}}$. Indeed, by the definition of $\nu$ we have that $\nu(\widetilde{\phi})=0$ if and only if $c_{+}(\widetilde{\phi})=c_{-}(\widetilde{\phi})=0$, which by Theorem 1.1 (iv) only implies that $\Pi(\widetilde{\phi})$ is the identity. On the other hand, the induced conjugation invariant pseudonorm on $\operatorname{Cont}_{0}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$, i.e. the pseudonorm $\nu_{\ast}$ defined by
@@
For every $\phi\in\operatorname{Cont}_{0}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ we have $\nu_{\ast}(\phi)\leq 2\pi+T_{\underline{w}}$.